### v2

# Prym loci of branched double coverings and generalized Andreotti-Mayer loci

Abstract.
The Andreotti-Mayer locus is a subset of the moduli space of principally polarized abelian varieties, defined by a condition on the dimension of the singular locus of the theta divisor. It is known that the Jacobian locus in the moduli space is an irreducible component of the Andreotti-Mayer locus. In this paper, we generalize the Andreotti-Mayer locus to the case of the moduli space of abelian varieties with non-principal polarization and prove that the Prym locus of branched double coverings is an irreducible component of the generalized Andreotti-Mayer locus.
2020 Mathematics Subject Classification:
14H40 (primary); 14H42, 14K25 (secondary).
The author is partially supported by JSPS KAKENHI Grant Number 20K03543.
1. Introduction
[…]
$\operatorname{Prym}_{g,2n}:\,\mathcal{R}_{g,2n}\longrightarrow\mathcal{A}_{d}^{\Delta};\,[\tilde{C}\overset{\phi}{\rightarrow}{C}]\longmapsto(P,[\mathcal{L}]),$
where $\mathcal{R}_{g,2n}$ denotes the moduli space of double coverings of curves of genus $g$ branched at $2n$ points, and $\mathcal{A}_{d}^{\Delta}$ denotes the moduli space of $d$-dimensional polarized abelian varieties of type $\Delta$. When $g=0$, the Prym variety $)\operatorname{Prym}_{0,2n}(\phi)=(J_{\tilde{C}},[\mathcal{O}_{J_{\tilde{C}}}(\Theta_{\tilde{C}}$)]) is the hyperelliptic Jacobian, and the Prym map $\operatorname{Prym}_{0,2n}$ is injective by the Torelli theorem. When $g=1$, it is proven in [8] that the Prym map $\operatorname{Prym}_{1,2n}$ is injective for $n\geq 3$. More generally, the following result holds:
Theorem 1.1 (Naranjo-Ortega [14]).
If $g\geq 1$ and $n\geq 3$, then the Prym map $\operatorname{Prym}_{g,2n}$ is injective.
Let $\mathcal{P}_{g,2n}\subset\mathcal{A}_{d}^{\Delta}$ be the closure of the image of the Prym map $\operatorname{Prym}_{g,2n}$. We aim to understand which kinds of polarized abelian varieties are contained in the Prym locus $\mathcal{P}_{g,2n}$. In the case where the polarization is principal, the characterization of the Jacobian locus has been classically studied as the Schottky problem. Let $\mathcal{J}_{d}\subset\mathcal{A}_{d}^{(1,\dots,1)}$ (resp. $\mathcal{H}_{d}\subset\mathcal{A}_{d}^{(1,\dots,1)}$) denote the closure of the set of points representing the Jacobian varieties of projective smooth curves (resp. hyperelliptic curves) of genus $d$. In this paper, we characterize $\mathcal{P}_{g,2n}$ in $\mathcal{A}_{d}^{\Delta}$, following the approach of Andreotti and Mayer [1] to characterize $\mathcal{J}_{d}$ and $\mathcal{H}_{d}$ in $\mathcal{A}_{d}^{(1,\dots,1)}$.
For an ample invertible sheaf $\mathcal{L}$ on an abelian variety $A$, we define the higher base locus of $\mathcal{L}$ by
$S(A,\mathcal{L})=\bigcap_{\Theta\in|\mathcal{L}|}\Theta_{\mathrm{sing}}\subset{A},$
where $\Theta_{\mathrm{sing}}$ denotes the singular locus of the hypersurface $\Theta\subset{A}$. If ample invertible sheaves $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ are algebraically equivalent, then $S(A,\mathcal{L}_{1})\simeq{S(A,\mathcal{L}_{2})}$, because $t_{x_{0}}^{*}\mathcal{L}_{1}\simeq\mathcal{L}_{2}$ by a translation
$t_{x_{0}}:A\longrightarrow{A};\,x\longmapsto{x+x_{0}}.$
For an integer $m\geq 0$, a generalization of the Andreotti-Mayer locus is defined by
$\mathcal{N}_{d,m}^{\Delta}=\{(A,[\mathcal{L}])\in\mathcal{A}_{d}^{\Delta}\mid\dim{S(A,\mathcal{L})}\geq{m}\}.$
When $\Delta=(1,\dots,1)$, the subset $\mathcal{N}_{d,m}^{(1,\dots,1)}\subset\mathcal{A}_{d}^{(1,\dots,1)}$ is the original Andreotti-Mayer locus, and the following results are known:
Theorem 1.2 (Andreotti-Mayer [1]).
  1. $(1)$
    If $d\geq 4$, then the Jacobian locus $\mathcal{J}_{d}$ is an irreducible component of $\mathcal{N}_{d,d-4}^{(1,\dots,1)}$.
  2. $(2)$
    If $d\geq 3$, then the hyperelliptic Jacobian locus $\mathcal{H}_{d}$ is an irreducible component of $\mathcal{N}_{d,d-3}^{(1,\dots,1)}$.
Theorem 1.3 (Debarre [4]).
If $d\geq 7$, then $\mathcal{P}_{d+1,0}$ is an irreducible component of $\mathcal{N}_{d,d-6}^{(1,\dots,1)}$, where $\mathcal{P}_{d+1,0}\subset\mathcal{A}_{d}^{(1,\dots,1)}$ denotes the Prym locus for unramified double coverings of curves of genus $d+1$.
In this paper, we prove the following:
Theorem 1.4.
If $n\geq 4$ and d=g+n1d=g+n-1, then the Prym locus $\mathcal{P}_{g,2n}\subset\mathcal{A}_{d}^{\Delta}$ is an irreducible component of $\mathcal{N}_{d,n-4}^{\Delta}$.
This is a generalization of Theorem 1.2 (2), because $\mathcal{P}_{0,2n}=\mathcal{H}_{d}\subset\mathcal{A}_{d}^{(1,\dots,1)}$ for d=n1d=n-1. According to the folk conjecture [7, Conjecture 3.15 (2)], it is expected that $\overline{\mathcal{N}_{d,d-3}^{(1,\dots,1)}\setminus\mathcal{N}_{d,d-2}^{(1,\dots,1)}}=\mathcal{H}_{d}$. More generally we expect that $\overline{\mathcal{N}_{d,n-4}^{\Delta}\setminus\mathcal{N}_{d,n-3}^{\Delta}}=\mathcal{P}_{g,2n}$ for d=g+n1d=g+n-1, but this remains an open problem.
This paper proceeds as follows: In Section 2, for the Prym variety $(P,[\mathcal{L}])$ of a branched double covering, we explicitly describe the higher base locus $S(P,\mathcal{L})\subset{P}$, and prove that $\mathcal{P}_{g,2n}\subset\mathcal{N}_{d,n-4}^{\Delta}$. In the proof, we provide the equation of the tangent cone $\mathcal{T}_{\Theta,x}\subset{T_{P,x}=T_{P,0}}$ of $\Theta\in|\mathcal{L}|$ at $x\in{S(P,\mathcal{L})}$. In Section 3, we prove that the equations of the tangent cones $\mathcal{T}_{\Theta,x}$ generate a certain subspace in $\operatorname{Sym}^{2}{T_{P,0}}^{\vee}$. In Section 4, following the approach in [1], we introduce a method to bound the dimension of an irreducible component $\mathcal{P}$ of $\mathcal{N}_{d,n-4}^{\Delta}$. When $\mathcal{P}_{g,2n}\subset\mathcal{P}$, by using independent equations of the tangent cones $\mathcal{T}_{\Theta,x}$, we can provide an upper bound for the dimension of $\mathcal{P}$. Combining this with the result from Section 3, we obtain $\dim{\mathcal{P}}\leq\dim{\mathcal{P}_{g,2n}}$, which implies that $\mathcal{P}=\mathcal{P}_{g,2n}$.
2. Higher base loci and the tangent cone of theta divisors
We assume that $n\geq 1$ and $g\geq 1$. Let $(P,[\mathcal{L}])$ be the Prym variety of $\tilde{C}\overset{\phi}{\rightarrow}{C}]\in\mathcal{R}_{g,2n}$. Following Mumford’s paper [11], we describe the base locus of the linear system $|\mathcal{L}|$. By [11, p333], we may assume that
[…]
and
$\psi^{*}\mathcal{O}_{J_{\tilde{C}}}(\Theta_{\tilde{C}})=\mathcal{O}_{J_{C}}(2\Theta_{C}),\quad\mathcal{L}=\mathcal{O}_{\tilde{C}}(\Theta_{\tilde{C}})|_{P},$
where $\psi:J_{C}\rightarrow{J_{\tilde{C}}}$ is the homomorphism defined by
$J_{C}=\operatorname{Pic}^{0}{(C)}\longrightarrow{J_{\tilde{C}}=\operatorname{Pic}^{0}{(\tilde{C})}};\,\xi\longmapsto\phi^{*}\xi.$
We define the homomorphism
$\pi:\,J_{C}\times{P}\longrightarrow{J_{\tilde{C}}};\,(\xi,\mathcal{F})\longmapsto{\mathcal{F}\otimes\phi^{*}\xi}.$
Then we have $\pi^{*}\mathcal{O}_{J_{\tilde{C}}}(\Theta_{\tilde{C}})={\mathrm{pr}_{1}^{*}\mathcal{O}_{J_{C}}(2\Theta_{C})\otimes{\mathrm{pr}_{2}^{*}\mathcal{L}}}$, where $\mathrm{pr}_{i}$ denotes the $i$-th projection of $J_{C}\times{P}$. Let $s\in{H^{0}(J_{\tilde{C}},\mathcal{O}_{J_{\tilde{C}}}(\Theta_{\tilde{C}}))}$ be a nontrivial section. Since $n\geq 1$, $\psi:J_{C}\rightarrow{J_{\tilde{C}}}$ is injective, and by [11, p335], we have
$\pi^{*}s=\sum_{i=1}^{2^{g}}\mathrm{pr}_{1}^{*}\alpha_{i}\otimes{\mathrm{pr}_{2}^{*}\beta_{i}}\in{H^{0}(J_{C}\times{P},\mathrm{pr}_{1}^{*}\mathcal{O}_{J_{C}}(2\Theta_{C})\otimes{\mathrm{pr}_{2}^{*}\mathcal{L}})}$
for some bases $\alpha_{1},\dots,\alpha_{2^{g}}\in{H^{0}(J_{C},\mathcal{O}_{J_{C}}(2\Theta_{C}))}$ and $\beta_{1},\dots,\beta_{2^{g}}\in{H^{0}(P,\mathcal{L})}$. This implies the following proposition for the base locus of $|\mathcal{L}|$. For $\mathcal{F}_{0}\in\operatorname{Pic}{(\tilde{C})}$, we denote the translation by
$t_{\mathcal{F}_{0}}:\,\operatorname{Pic}{(\tilde{C})}\longrightarrow\operatorname{Pic}{(\tilde{C})};\ \mathcal{F}\longmapsto{\mathcal{F}\otimes{\mathcal{F}_{0}}}.$
Proposition 2.1 ([11] p334).
The base locus of the linear system $|\mathcal{L}|$ is
$\operatorname{Bs}{|\mathcal{L}|}=\{\mathcal{F}\in{P}\mid{\psi(J_{C})\subset t_{\mathcal{F}}^{-1}(\Theta_{\tilde{C}})}\}=\bigcap_{\xi\in\operatorname{Pic}^{0}{(C)}}t_{\phi^{*}\xi}^{-1}(\Theta_{\tilde{C}})|_{P}.$
More precisely, $H^{0}(P,\mathcal{L})$ is generated by defining sections of the hypersurfaces $t_{\phi^{*}\xi}^{-1}(\Theta_{\tilde{C}})|_{P}\subset{P}$ for $\xi\in\operatorname{Pic}^{0}{(C)}$.
For $\tilde{C}\overset{\phi}{\rightarrow}{C}]\in\mathcal{R}_{g,2n}$, there exists a unique invertible sheaf $\eta_{\phi}\in\operatorname{Pic}^{n}{(C)}$ such that $\Omega_{\tilde{C}}^{1}=\phi^{*}(\Omega_{C}^{1}\otimes\eta_{\phi})$. Let
$P_{\phi}=\operatorname{Nm}^{-1}(\Omega_{C}^{1}\otimes\eta_{\phi})\subset\operatorname{Pic}^{\tilde{g}-1}{(\tilde{C})}$
be the fiber of the norm map at $\Omega_{C}^{1}\otimes\eta_{\phi}\in\operatorname{Pic}{(C)}$, where $\tilde{g}=2g+n-1$ is the genus of $\tilde{C}$. We set
$W_{\tilde{C}}=\{\mathcal{F}\in\operatorname{Pic}^{\tilde{g}-1}{(\tilde{C})}\mid h^{0}(\tilde{C},\mathcal{F})\geq 1\}$
and $\mathcal{L}_{\phi}=\mathcal{O}_{\operatorname{Pic}^{\tilde{g}-1}{(\tilde{C})}}(W_{\tilde{C}})|_{P_{\phi}}$. Then there exists $\mathcal{F}_{0}\in{P_{\phi}}$ such that
[…]
hence the translation $t_{\mathcal{F}_{0}}$ induces isomorphisms $P=\operatorname{Ker}{(\operatorname{Nm})}\overset{\simeq}{\rightarrow}P_{\phi}$ and $\operatorname{Bs}{|\mathcal{L}|}\overset{\simeq}{\rightarrow}\operatorname{Bs}{|\mathcal{L}_{\phi}|}$. In [13], Naranjo and Ortega gave an explicit description for $\operatorname{Bs}{|\mathcal{L}_{\phi}|}$. For $j\geq 0$, we denote by $\tilde{C}^{(j)}$ the $j$-th symmetric product of $\tilde{C}$, which parameterizes effective divisors of degree $j$ on $\tilde{C}$. We define the subvariety in $\operatorname{Pic}^{\tilde{g}-1}{(\tilde{C})}$ by
$B_{i}=\operatorname{Image}{(\tilde{C}^{(n-2i)}\times\operatorname{Pic}^{g-1+i}{(C)}\longrightarrow\operatorname{Pic}^{\tilde{g}-1}{(\tilde{C})};\,(D,\mathcal{G})\longmapsto\mathcal{O}_{\tilde{C}}(D)\otimes\phi^{*}\mathcal{G})}$
for $2i\leq{n}$. When $2i>n$, we set $B_{i}=\emptyset$.
Proposition 2.2 ([13] Proposition 1.6.).
${\operatorname{Bs}{|\mathcal{L}_{\phi}|}=B_{1}\cap{P_{\phi}}}$.
When we set $W_{\xi}=t_{\phi^{*}\xi}^{-1}(W_{\tilde{C}})|_{P_{\phi}}\in|\mathcal{L}_{\phi}|$ for $\xi\in\operatorname{Pic}^{0}{(C)}$, as in Proposition 2.1 we have
${\operatorname{Bs}{|\mathcal{L}_{\phi}|}=\bigcap_{\xi\in\operatorname{Pic}^{0}{(C)}}W_{\xi}=B_{1}\cap{P_{\phi}}},$
and $H^{0}(P_{\phi},\mathcal{L}_{\phi})$ is generated by defining sections of $W_{\xi}$ for $\xi\in\operatorname{Pic}^{0}{(C)}$. We define $S_{i}(P_{\phi},\mathcal{L}_{\phi})=\bigcap_{\Theta\in|\mathcal{L}_{\phi}|}\Theta_{\geq{i}}$, where $\Theta_{\geq{i}}$ denotes the higher multiplicity locus
$\Theta_{\geq{i}}=\{\mathcal{F}\in\Theta\mid\operatorname{mult}_{\mathcal{F}}{\Theta}\geq{i}\}\subset{P_{\phi}}.$
By the next proposition, we have
$S(P,\mathcal{L})\overset{\simeq}{\longrightarrow}S_{2}(P_{\phi},\mathcal{L}_{\phi})=\bigcap_{\xi\in\operatorname{Pic}^{0}{(C)}}W_{\xi,\mathrm{sing}}=B_{2}\cap{P_{\phi}}.$
Proposition 2.3.
When $i\geq 1$,
$S_{i}(P_{\phi},\mathcal{L}_{\phi})=\bigcap_{\xi\in\operatorname{Pic}^{0}{(C)}}W_{\xi,\geq{i}}=B_{i}\cap{P_{\phi}}.$
We prepare some lemmas to prove Proposition 2.3.
Lemma 2.4.
If $\mathcal{F}\in{B_{i}\cap{P_{\phi}}}$, then $\operatorname{mult}_{\mathcal{F}}{W_{\xi}}\geq{i}$ for any $\xi\in\operatorname{Pic}^{0}{(C)}$.
Proof.
There exist $D\in\tilde{C}^{(n-2i)}$ and $\mathcal{G}\in\operatorname{Pic}^{g-1+i}{(\mathcal{G})}$ such that $\mathcal{F}=\mathcal{O}_{\tilde{C}}(D)\otimes\phi^{*}\mathcal{G}$. Since
$\mathcal{G}\otimes\xi\subset\phi_{*}\phi^{*}(\mathcal{G}\otimes\xi)\subset\phi_{*}(\mathcal{O}_{\tilde{C}}(D)\otimes\phi^{*}(\mathcal{G}\otimes\xi))=\phi_{*}(\mathcal{F}\otimes\phi^{*}\xi),$
by the Riemann’s singularity theorem, we have
$\operatorname{mult}_{\mathcal{F}}{W_{\xi}}\geq\operatorname{mult}_{\mathcal{F}}{t_{\phi^{*}\xi}^{-1}(W_{\tilde{C}})}=h^{0}(\tilde{C},\mathcal{F}\otimes\phi^{*}\xi)\geq{h^{0}(C,\mathcal{G}\otimes\xi)}\geq{i}.$
The following Mumford’s exact sequence is essential for the proof of Proposition 2.3.
Lemma 2.5 ([11] p338).
Let $D$ be an effective divisor on $\tilde{C}$. If $\phi^{*}p\nleq{D}$ for any $p\in{C}$, then there exists an exact sequence
$0\longrightarrow\mathcal{O}_{C}\longrightarrow\phi_{*}\mathcal{O}_{\tilde{C}}(D)\longrightarrow\operatorname{Nm}{(\mathcal{O}_{\tilde{C}}(D))}\otimes\eta_{\phi}^{\vee}\longrightarrow 0$
of locally free sheaves on $C$.
In the following, we write $H^{i}(\mathcal{G})=H^{i}(C,\mathcal{G})$ for any sheaf $\mathcal{G}$ on $C$. When $0<2i\leq{n}$, for $E\in{C^{(n-2i)}}$ and $\eta\in\operatorname{Pic}^{n}{(C)}$, we define
$K_{i}(E,\eta)=\{\mathcal{G}\in\operatorname{Pic}^{g-1+i}{(C)}\mid h^{0}(\mathcal{G})=h^{0}(\Omega_{C}^{1}\otimes\eta\otimes\mathcal{O}_{C}(-E)\otimes\mathcal{G}^{\vee})=i\}.$
For $\mathcal{G}\in{K_{i}(E,\eta)}$, we define the hypersurface $\mathcal{V}_{E,\mathcal{G}}\subset{H^{0}(\Omega_{C}^{1}\otimes\eta)^{\vee}}$ of degree $i$ by
$\det{\begin{pmatrix}\mu(s_{1},t_{1})&&\mu(s_{1},t_{i})\\ &\cdots&\\ \mu(s_{i},t_{1})&&\mu(s_{i},t_{i})\\ \end{pmatrix}\in\operatorname{Sym}^{i}{H^{0}(\Omega_{C}^{1}\otimes\eta)},}$
where $s_{1},\dots,s_{i}\in{H^{0}(\mathcal{G})}$ and $t_{1},\dots,t_{i}\in{H^{0}(\Omega_{C}^{1}\otimes\eta\otimes\mathcal{O}_{C}(-E)\otimes\mathcal{G}^{\vee})}$ are bases, and
$\mu:\,H^{0}(\mathcal{G})\times H^{0}(\Omega_{C}^{1}\otimes\eta\otimes\mathcal{O}_{C}(-E)\otimes\mathcal{G}^{\vee})\longrightarrow H^{0}(\Omega_{C}^{1}\otimes\eta)$
denotes the composition of the multiplication map and the injective homomorphism $H^{0}(\Omega_{C}^{1}\otimes\eta\otimes\mathcal{O}_{C}(-E))\hookrightarrow{H^{0}(\Omega_{C}^{1}\otimes\eta)}$. We note that $\mathcal{V}_{E,\mathcal{G}}$ does not depend on the choice of bases $s_{1},\dots,s_{i}$ and $t_{1},\dots,t_{i}$.
Lemma 2.6.
Let $D\in\tilde{C}^{(n-2i)}$ and $\mathcal{G}\in\operatorname{Pic}^{g-1+i}{(C)}$ be such that $\mathcal{F}=\mathcal{O}_{\tilde{C}}(D)\otimes\phi^{*}\mathcal{G}\in{P_{\phi}}$. For $\xi\in\operatorname{Pic}^{0}{(C)}$, the following conditions are equivalent:
  1. $(1)$
    $\operatorname{mult}_{\mathcal{F}}{W_{\xi}}=i$.
  2. $(2)$
    $\mathcal{G}\otimes\xi\in{K_{i}(\phi(D),\eta_{\phi})}$, and $\phi^{*}p\nleq{D}$ for any $p\in{C}$.
If these conditions are satisfied, then the tangent cone $\mathcal{T}_{W_{\xi},\mathcal{F}}$ of $W_{\xi}$ at $\mathcal{F}$ is
$\mathcal{V}_{\phi(D),\mathcal{G}\otimes\xi}\subset{H^{0}(\Omega_{C}^{1}\otimes\eta_{\phi})^{\vee}}\simeq{T_{P_{\phi},\mathcal{F}}}.$
Proof.
First, assume condition (1). If $\phi^{*}p\leq{D}$ for some $p\in{C}$, then $\mathcal{F}\in{B_{i+1}\cap{P_{\phi}}}$, and by Lemma 2.4, we have $\operatorname{mult}_{\mathcal{F}}{W_{\xi}}\geq{i+1}$, which contradicts the assumption that $\operatorname{mult}_{\mathcal{F}}{W_{\xi}}=i$. Thus, $\phi^{*}p\nleq{D}$ for any $p\in{C}$. Since $\operatorname{Nm}{(\mathcal{O}_{\tilde{C}}(D))}\otimes\mathcal{G}^{\otimes 2}=\Omega_{C}^{1}\otimes\eta_{\phi}$, by Lemma 2.5, we obtain the exact sequence
$0\longrightarrow\mathcal{G}\otimes\xi\longrightarrow\phi_{*}(\mathcal{F}\otimes\phi^{*}\xi)\longrightarrow\Omega_{C}^{1}\otimes\mathcal{G}^{\vee}\otimes\xi\longrightarrow 0.$
Since
$i=\operatorname{mult}_{\mathcal{F}}{W_{\xi}}\geq\operatorname{mult}_{\mathcal{F}}{t_{\phi^{*}\xi}^{-1}(W_{\tilde{C}})}=h^{0}(\tilde{C},\mathcal{F}\otimes\phi^{*}\xi)\geq h^{0}(\mathcal{G}\otimes\xi)\geq{i},$
it follows that $h^{0}(\mathcal{G}\otimes\xi)=i$ and $h^{1}(\mathcal{G}\otimes\xi)=0$. Therefore, $\mathcal{G}\otimes\xi\in{K_{i}(\phi(D),\eta_{\phi})}$, because
$h^{0}(\Omega_{C}^{1}\otimes\eta_{\phi}\otimes\operatorname{Nm}{(\mathcal{O}_{\tilde{C}}(-D))}\otimes(\mathcal{G}\otimes\xi)^{\vee})=h^{0}(\mathcal{G}\otimes\xi^{\vee})$
$=$ $h^{0}(\Omega_{C}^{1}\otimes\mathcal{G}^{\vee}\otimes\xi)+i=h^{0}(\tilde{C},\mathcal{F}\otimes\phi^{*}\xi)-h^{0}(\mathcal{G}\otimes\xi)+i=i.$
Next, assume condition (2). By using Lemma 2.5, we obtain
$\operatorname{mult}_{\mathcal{F}}{t_{\phi^{*}\xi}^{-1}(W_{\tilde{C}})}=h^{0}(\tilde{C},\mathcal{F}\otimes\phi^{*}\xi)=h^{0}(\mathcal{G}\otimes\xi)=i.$
Let $\sigma:\tilde{C}\rightarrow\tilde{C}$ be the covering involution of $\phi$. Since $\mathcal{F}\in{P_{\phi}}$, we have
$\Omega_{\tilde{C}}^{1}=\phi^{*}(\Omega_{C}^{1}\otimes\eta_{\phi})=\mathcal{F}\otimes\sigma^{*}\mathcal{F}=\mathcal{F}\otimes\mathcal{O}_{\tilde{C}}(\sigma^{*}D)\otimes\phi^{*}\mathcal{G}.$
The injective homomorphisms
$\mathcal{G}\otimes\xi\hookrightarrow\phi_{*}(\mathcal{O}_{\tilde{C}}(D))\otimes\mathcal{G}\otimes\xi=\phi_{*}(\mathcal{O}_{\tilde{C}}(D)\otimes\phi^{*}(\mathcal{G}\otimes\xi))=\phi_{*}(\mathcal{F}\otimes\phi^{*}\xi)$
and
$\Omega_{C}^{1}\otimes\eta_{\phi}\otimes\operatorname{Nm}{(\mathcal{O}_{\tilde{C}}(-D))}\otimes(\mathcal{G}\otimes\xi)^{\vee}=\mathcal{G}\otimes\xi^{\vee}\\ \hookrightarrow\phi_{*}(\mathcal{O}_{\tilde{C}}(\sigma^{*}D))\otimes\mathcal{G}\otimes\xi^{\vee}=\phi_{*}(\mathcal{O}_{\tilde{C}}(\sigma^{*}D)\otimes\phi^{*}(\mathcal{G}\otimes\xi^{\vee}))=\phi_{*}(\Omega^{1}_{\tilde{C}}\otimes(\mathcal{F}\otimes\phi^{*}\xi)^{\vee})$
induce the isomorphisms
$H^{0}(C,\mathcal{G}\otimes\xi)\overset{\simeq}{\longrightarrow}H^{0}(\tilde{C},\mathcal{F}\otimes\phi^{*}\xi)$
and
$H^{0}(C,\Omega_{C}^{1}\otimes\eta_{\phi}\otimes\operatorname{Nm}{(\mathcal{O}_{\tilde{C}}(-D))}\otimes(\mathcal{G}\otimes\xi)^{\vee})\overset{\simeq}{\longrightarrow}H^{0}(\tilde{C},\Omega^{1}_{\tilde{C}}\otimes(\mathcal{F}\otimes\phi^{*}\xi)^{\vee}).$
From the commutative diagram
[…]
$(j^{\vee})^{-1}(\mathcal{V}_{\phi(D),\mathcal{G}\otimes\xi})\subsetneq{H^{0}(\tilde{C},\Omega_{\tilde{C}}^{1})^{\vee}}\simeq{T_{\operatorname{Pic}^{\tilde{g}-1}{(\tilde{C})},\mathcal{F}}}$,
where $j^{\vee}:\,H^{0}(\tilde{C},\Omega_{\tilde{C}}^{1})^{\vee}\rightarrow{H^{0}(\Omega_{C}^{1}\otimes\eta_{\phi})^{\vee}}$ is the dual of the injective homomorphism
$j:\,H^{0}(\Omega_{C}^{1}\otimes\eta_{\phi})\longrightarrow H^{0}(\tilde{C},\phi^{*}(\Omega_{C}^{1}\otimes\eta_{\phi}))=H^{0}(\tilde{C},\Omega_{\tilde{C}}^{1}).$
Since $T_{P_{\phi},\mathcal{F}}\subset{H^{0}(\tilde{C},\Omega_{\tilde{C}}^{1})^{\vee}}$ is isomorphic to $H^{0}(\Omega_{C}^{1}\otimes\eta_{\phi})^{\vee}$ via $j^{\vee}$, the tangent cone of $W_{\xi}$ at $\mathcal{F}$ is
$T_{P_{\phi},\mathcal{F}}\cap(j^{\vee})^{-1}(\mathcal{V}_{\phi(D),\mathcal{G}\otimes\xi})\simeq\mathcal{V}_{\phi(D),\mathcal{G}\otimes\xi}\subsetneq H^{0}(\Omega_{C}^{1}\otimes\eta_{\phi})^{\vee},$
and hence we conclude that $\operatorname{mult}_{\mathcal{F}}{W_{\xi}}=i$. ∎
Proof of Proposition 2.3.
Let $\Theta\in|\mathcal{L}_{\phi}|$ be the hypersurface defined by a section $\beta\in{H^{0}(P_{\phi},\mathcal{L}_{\phi})}$. We have $\Theta_{\geq{i}}\supset\bigcap_{\xi\in\operatorname{Pic}^{0}{(C)}}W_{\xi,\geq{i}}$, because $\beta$ is a linear combination of defining sections of $W_{\xi}$ for $\xi\in\operatorname{Pic}^{0}{(C)}$. Hence, we have $S_{i}(P_{\phi},\mathcal{L}_{\phi})=\bigcap_{\xi\in\operatorname{Pic}^{0}{(C)}}W_{\xi,\geq{i}}$.
By Lemma 2.4, we have $B_{i}\cap{P_{\phi}}\subset\bigcap_{\xi\in\operatorname{Pic}^{0}{(C)}}W_{\xi,\geq{i}}.$ By induction on $i$, we prove $B_{i}\cap{P_{\phi}}=\bigcap_{\xi\in\operatorname{Pic}^{0}{(C)}}W_{\xi,\geq{i}}$. When $i=1$, this is true by Proposition 2.2. We assume $i\geq 2$. By the induction assumption,
$\bigcap_{\xi\in\operatorname{Pic}^{0}{(C)}}W_{\xi,\geq{i}}\subset\bigcap_{\xi\in\operatorname{Pic}^{0}{(C)}}W_{\xi,\geq{i-1}}=B_{i-1}\cap{P_{\phi}}.$
When n<2(i1)n<2(i-1), we have defined that $B_{i-1}=B_{i}=\emptyset$, and it follows that $\bigcap_{\xi\in\operatorname{Pic}^{0}{(C)}}W_{\xi,\geq{i}}=\emptyset$ and $B_{i}\cap{P_{\phi}}=\emptyset$. When $n\geq 2(i-1)$, for $\mathcal{F}\in\bigcap_{\xi\in\operatorname{Pic}^{0}{(C)}}W_{\xi,\geq{i}}$, there exist $D\in\tilde{C}^{(n-2(i-1))}$ and $\mathcal{G}\in\operatorname{Pic}^{g-1+(i-1)}{(C)}$ such that
$\mathcal{F}=\mathcal{O}_{\tilde{C}}(D)\otimes\phi^{*}\mathcal{G},\quad\Omega_{C}^{1}\otimes\eta_{\phi}=\operatorname{Nm}{(\mathcal{O}_{\tilde{C}}(D))}\otimes\mathcal{G}^{\otimes 2}.$
Since $\deg{\mathcal{G}}=g+i-2$, we can find $\xi_{0}\in\operatorname{Pic}^{0}{(C)}$ such that $h^{0}(\mathcal{G}\otimes\xi_{0})=i-1$ and $h^{0}(\mathcal{G}\otimes\xi_{0}^{\vee})=i-1$. Then, we have $\mathcal{G}\otimes\xi_{0}\in{K_{i-1}(\phi(D),\eta_{\phi})}$. If $\mathcal{F}\notin{B_{i}}$, then $\phi^{*}p\nleq{D}$ for any $p\in{D}$, and by Lemma 2.6, we have $\operatorname{mult}_{\mathcal{F}}{W_{\xi_{0}}}=i-1$. This is a contradiction to $\mathcal{F}\in\bigcap_{\xi\in\operatorname{Pic}^{0}{(C)}}{W_{\xi,\geq{i}}}$, hence $\mathcal{F}\in{B_{i}}$. ∎
Proposition 2.7.
$\dim{(B_{i}\cap{P_{\phi}})}=n-2i$ for $0<2i\leq{n}$. If $0<2i<n$, then $B_{i}\cap{P_{\phi}}$ is irreducible. If $0<2i=n$, then ${(B_{i}\cap{P_{\phi}})}=4^{g}$.
Proof.
When $0<2i\leq{n}$, we define
$Y_{i}=\{(D,\mathcal{G})\in\tilde{C}^{(n-2i)}\times\operatorname{Pic}^{g-1+i}{(C)}\mid\operatorname{Nm}{(\mathcal{O}_{\tilde{C}}(D))\otimes\mathcal{G}^{\otimes 2}}=\Omega_{C}^{1}\otimes\eta_{\phi}\}.$
The projection
$g_{i}:\,Y_{i}\longrightarrow\tilde{C}^{(n-2i)};\,(D,\mathcal{G})\longmapsto{D}$
is a finite étale covering of degree $4^{g}$, hence $Y_{i}$ is smooth. When $n=2i$,
$Y_{i}\longrightarrow{B_{i}\cap{P_{\phi}}};\,(0,\mathcal{G})\longmapsto\phi^{*}\mathcal{G}$
is bijective, hence ${(B_{i}\cap{P_{\phi}})}=\#{Y_{i}}=4^{g}$. Assume $n\geq 2i+1$. Let $\gamma:[0,1]\rightarrow\tilde{C}$ be a loop on $\tilde{C}$, and let $\iota$ be the morphism defined by
$\iota:\,C\longrightarrow{J_{C}=\operatorname{Pic}^{0}{(C)}};\,p\longmapsto\mathcal{O}_{C}(p-\phi(\gamma(0))).$
Then, there is a unique lift $\delta:[0,1]\rightarrow{H^{0}(\Omega_{C}^{1})^{\vee}}$ of $\iota\circ\phi\circ\gamma$ such that $\delta(0)=0$:
[…]
For $t\in[0,1]$, we denote by $\xi(t)\in\operatorname{Pic}^{0}{(C)}$ the invertible sheaf corresponding to $\bigl[\frac{1}{2}\delta(t)\bigr]\in{H^{0}(\Omega_{C}^{1})^{\vee}/H_{1}(C,\mathbb{Z})}=J_{C}$. If $(D,\mathcal{G})=(\gamma(0)+D^{\prime},\mathcal{G})\in{Y_{i}}$, then the path
$0,1]\longrightarrow{Y_{i}};\,t\longmapsto(\gamma(t)+D^{\prime},\mathcal{G}\otimes\xi(t)^{\vee})$
on $Y_{i}$ is the lift of the loop
$0,1]\longrightarrow\tilde{C}^{(n-2i)};\,t\longmapsto\gamma(t)+D^{\prime}$
on $\tilde{C}^{(n-2i)}$. Since $\phi_{*}:H_{1}(\tilde{C},\mathbb{Z})\rightarrow{H_{1}(C,\mathbb{Z})}$ is surjective,
$H_{1}(\tilde{C},\mathbb{Z})\longrightarrow{J_{C,2}=\operatorname{Ker}{([2]_{J_{C}})}};\,[\gamma]\longmapsto\xi(1)=\Bigl[\frac{1}{2}\delta(1)\Bigr]$
is surjective. This means that $\pi_{1}(\tilde{C}^{(n-2i)},D)$ acts transitively on the fiber $g_{i}^{-1}(D)$, and it follows that $Y_{i}$ is connected. Since $Y_{i}$ is also smooth, it is irreducible. Let $U_{i}$ be the subset of $\tilde{C}^{(n-2i)}$ defined by
$U_{i}=\{D\in\tilde{C}^{(n-2i)}\mid h^{0}(\tilde{C},\mathcal{O}_{\tilde{C}}(D))=1,\,\text{$\phi^{*}p\nleq{D}$ for any $p\in{C}$}\}.$
Since $0<n-2i\leq 2g+n-1=\tilde{g}$, it is a non-empty open subset of $\tilde{C}^{(n-2i)}$. For $(D_{1},\mathcal{G}_{1}),(D_{2},\mathcal{G}_{2})\in g_{i}^{-1}(U_{i})$, we have
$h^{0}(\operatorname{Nm}{(\mathcal{O}_{\tilde{C}}(D_{1}))}\otimes\eta_{\phi}^{\vee}\otimes\mathcal{G}_{1}\otimes\mathcal{G}_{2}^{\vee})=h^{0}(\Omega_{C}^{1}\otimes\mathcal{G}_{1}^{\vee}\otimes\mathcal{G}_{2}^{\vee})=0,$
because $\deg{(\Omega_{C}^{1}\otimes\mathcal{G}_{1}^{\vee}\otimes\mathcal{G}_{2}^{\vee})}=-2i<0$. If $\mathcal{O}_{\tilde{C}}(D_{1})\otimes\phi^{*}\mathcal{G}_{1}=\mathcal{O}_{\tilde{C}}(D_{2})\otimes\phi^{*}\mathcal{G}_{2}$, then by Lemma 2.5, we have
$h^{0}(\mathcal{G}_{1}\otimes\mathcal{G}_{2}^{\vee})=h^{0}(\tilde{C},\mathcal{O}_{\tilde{C}}(D_{1})\otimes\phi^{*}(\mathcal{G}_{1}\otimes\mathcal{G}_{2}^{\vee}))=h^{0}(\tilde{C},\mathcal{O}_{\tilde{C}}(D_{2}))=1,$
hence $\mathcal{G}_{1}\otimes\mathcal{G}_{2}^{\vee}=\mathcal{O}_{C}$ and $D_{1}=D_{2}$. This means that the morphism
$Y_{i}\longrightarrow{B_{i}\cap{P}_{\phi}};\,(D,\mathcal{G})\longmapsto\mathcal{O}_{\tilde{C}}(D)\otimes\phi^{*}\mathcal{G}$
is birational. Hence, $B_{i}\cap{P_{\phi}}$ is irreducible, and
$\dim{(B_{i}\cap{P}_{\phi})}=\dim{Y_{i}}=n-2i.$
By Proposition 2.3 and Proposition 2.7, we have the following:
Corollary 2.8.
Let $(P,[\mathcal{L}])$ be the Prym variety of $\tilde{C}\overset{\phi}{\rightarrow}{C}]\in\mathcal{R}_{g,2n}$ for $n\geq 1$ and $g\geq 1$.
  1. $(1)$
    1. (a)
      $\dim{\operatorname{Bs}{|\mathcal{L}|}}=n-2$ and $\operatorname{Bs}{|\mathcal{L}|}$ is irreducible for $n\geq 3$,
    2. (b)
      $\dim{\operatorname{Bs}{|\mathcal{L}|}}=0$ and ${\operatorname{Bs}{|\mathcal{L}|}}=4^{g}$ for $n=2$,
    3. (c)
      $\operatorname{Bs}{|\mathcal{L}|}=\emptyset$ for $n=1$,
  2. $(2)$
    1. (a)
      $\dim{S(P,\mathcal{L})}=n-4$ and $S(P,\mathcal{L})$ is irreducible for $n\geq 5$,
    2. (b)
      $\dim{S(P,\mathcal{L})}=0$ and ${S(P,\mathcal{L})}=4^{g}$ for $n=4$,
    3. (c)
      $S(P,\mathcal{L})=\emptyset$ for $1\leq{n}\leq 3$.
3. Quadric hypersurfaces containing projective curves
Let $C$ be a projective smooth curve of genus $g$, and let $\eta$ be an invertible sheaf on $C$ of degree $n\geq 3$. By [10, p55, Corollary], the invertible sheaf $\Omega_{C}^{1}\otimes\eta$ is very ample, and the multiplication map
$\rho_{i}:\,\operatorname{Sym}^{i}{H^{0}(\Omega_{C}^{1}\otimes\eta)}\longrightarrow H^{0}((\Omega_{C}^{1}\otimes\eta)^{\otimes{i}})$
is surjective for $i\geq 1$. When $0<{2i}\leq{n}$, we have defined the hypersurface $\mathcal{V}_{E,\mathcal{G}}\subset{H^{0}(\Omega_{C}^{1}\otimes\eta)^{\vee}}=\operatorname{Spec}{(\operatorname{Sym}{H^{0}(\Omega_{C}^{1}\otimes\eta)})}$ for $E\in{C^{(n-2i)}}$ and $\mathcal{G}\in{K_{i}(E,\eta)}$. Let $Q_{E,\mathcal{G}}\subset{\mathbb{P}(H^{0}(\Omega_{C}^{1}\otimes\eta))}=\operatorname{Proj}{(\operatorname{Sym}{H^{0}(\Omega_{C}^{1}\otimes\eta)})}$ be the projectivization of $\mathcal{V}_{E,\mathcal{G}}$. Then we have
$Q_{E,\mathcal{G}}=\bigcup_{D\in|\mathcal{G}|}\mathbb{P}(H^{0}(\Omega_{C}^{1}\otimes\eta)/V_{E+D})=\bigcup_{D\in|\Omega_{C}^{1}\otimes\eta\otimes\mathcal{O}_{C}(-E)\otimes\mathcal{G}^{\vee}|}\mathbb{P}(H^{0}(\Omega_{C}^{1}\otimes\eta)/V_{E+D})$
in $\mathbb{P}(H^{0}(\Omega_{C}^{1}\otimes\eta))$, where
$V_{E+D}=\operatorname{Image}{(H^{0}(\Omega_{C}^{1}\otimes\eta\otimes\mathcal{O}_{C}(-E-D))\hookrightarrow{H^{0}(\Omega_{C}^{1}\otimes\eta)})}.$
Let
$\Phi=\Phi_{|\Omega_{C}^{1}\otimes\eta|}:\,C\longrightarrow\mathbb{P}(H^{0}(\Omega_{C}^{1}\otimes\eta))=\mathbb{P}^{d-1}$
be the closed immersion defined by the very ample linear system $|\Omega_{C}^{1}\otimes\eta|$. When $i\geq 2$, the defining section of $Q_{E,\mathcal{G}}\subset{\mathbb{P}^{d-1}}$ is contained in
$\operatorname{Ker}{(\rho_{i})}\subset\operatorname{Sym}^{i}{H^{0}(\Omega_{C}^{1}\otimes\eta)}={H^{0}(\mathbb{P}^{d-1},\mathcal{O}_{\mathbb{P}^{d-1}}(i))},$
hence, by the commutative diagram
[…]
we have $\Phi(C)\subset{Q_{E,\mathcal{G}}}$.
Proposition 3.1.
Let $\Sigma\subset{C}$ be a finite subset. If $n\geq 4$ and $C$ is an elliptic curve or a hyperelliptic curve, then $\operatorname{Ker}{(\rho_{2})}$ is generated by the defining quadratics of $Q_{E,\mathcal{G}}$ for $E\in{(C\setminus\Sigma)^{(n-4)}}$ and $\mathcal{G}\in{K_{2}(E,\eta)}$.
First we prove it in the case where $n=4$.
Lemma 3.2.
If $n=4$ and $C$ is an elliptic curve or a hyperelliptic curve, then $\operatorname{Ker}{(\rho_{2})}$ is generated by the defining quadratics of $Q_{0,\mathcal{G}}$ for $\mathcal{G}\in{K_{2}(0,\eta)}$.
Proof.
Since $C$ is an elliptic curve or a hyperelliptic curve, there exists an invertible sheaf $\mathcal{H}\in\operatorname{Pic}^{2}{(C)}$ such that $h^{0}(\mathcal{H})=2$. Let $U\subset{C^{g+1}}$ be the set of $(p_{1},\dots,p_{g+1})\in{C^{g+1}}$ satisfying the following conditions:
  1. $(1)$
    $p_{i}\neq{p_{j}}$ for $1\leq{i}<j\leq{g+1}$,
  2. $(2)$
    $\mathcal{O}_{C}(p_{1}+\dots+p_{g+1})\in{K_{2}(0,\eta)}$,
  3. $(3)$
    $\mathcal{H}\otimes\mathcal{O}_{C}(p_{1}+\dots+p_{g+1}-p_{i}-p_{j})\in{K_{2}(0,\eta)}$ for $1\leq{i}<j\leq{g+1}$,
  4. $(4)$
    $h^{0}(\mathcal{H}^{\vee}\otimes\mathcal{O}_{C}(p_{1}+\dots+p_{g+1}))=0$,
  5. $(5)$
    $h^{0}(\Omega_{C}^{1}\otimes\eta\otimes\mathcal{H}^{\vee}\otimes\mathcal{O}_{C}(-p_{1}-\dots-p_{g+1}))=0$.
For $\mathcal{G}\in\operatorname{Pic}^{g+1}{(C)}$, the condition $\mathcal{G}\in{K(0,\eta)}$ is equivalent to
$h^{0}(\Omega_{C}^{1}\otimes\mathcal{G}^{\vee})=h^{0}(\mathcal{G}\otimes\eta^{\vee})=0.$
Since
$\deg{(\Omega_{C}^{1}\otimes\mathcal{G}^{\vee})}=\deg{(\mathcal{G}\otimes\eta^{\vee})}=g-3<g,$
$K_{2}(0,\eta)=\operatorname{Pic}^{g+1}{(C)}$ for $g\leq 2$, and $K_{2}(0,\eta)$ is a non-empty open subset of $\operatorname{Pic}^{g+1}{(C)}$ with $\dim{(\operatorname{Pic}^{g+1}{(C)}\setminus{K_{2}(0,\eta)})}=g-3$ for $g\geq 3$. Hence the set of $(p_{1},\dots,p_{g+1})\in{C^{g+1}}$ satisfying (2) is a non-empty open subset of $C^{g+1}$. Since the image of the map
$C^{g+1}\longrightarrow\operatorname{Pic}^{g+1}{(C)};\,(p_{1},\dots,p_{g+1})\longmapsto\mathcal{H}\otimes\mathcal{O}_{C}(p_{1}+\dots+p_{g+1}-p_{i}-p_{j})$
has dimension g1g-1 for $1\leq{i}<j\leq{g+1}$, the set of $(p_{1},\dots,p_{g+1})\in{C^{g+1}}$ satisfying (3) is a non-empty open subset of $C^{g+1}$. Since (1), (4) and (5) are also open conditions, $U$ is a non-empty open subset of $C^{g+1}$. We fix $(p_{1},\dots,p_{g+1})\in{U}$, and set $\mathcal{G}_{0}=\mathcal{O}_{C}(p_{1}+\dots+p_{g+1})$ and $\mathcal{G}_{ij}=\mathcal{H}\otimes\mathcal{O}_{C}(p_{1}+\dots+p_{g+1}-p_{i}-p_{j})$ for $1\leq{i}<j\leq{g+1}$. We show that the defining quadratics of $Q_{0,\mathcal{G}}$ and $Q_{0,\mathcal{G}_{ij}}$ form a basis of $\operatorname{Ker}{(\rho_{2})}$. For $E_{0}\neq{E_{\infty}}\in|\mathcal{H}|$, there is a rational function $f$ on $C$ such that $\operatorname{Div}{(f)}=E_{0}-E_{\infty}$. Let $\alpha_{ij}\in{H^{0}(\mathcal{G}_{ij})}$ be a section defining the divisor $E_{\infty}+p_{1}+\dots+p_{g+1}-p_{i}-p_{j}\in|\mathcal{G}_{ij}|$. Then $\alpha^{\prime}_{ij}=f\alpha_{ij}\in{H^{0}(\mathcal{G}_{ij})}$ defines the divisor $E_{0}+p_{1}+\dots+p_{g+1}-p_{i}-p_{j}\in|\mathcal{G}_{ij}|$, and by (3), $\alpha_{ij},\alpha^{\prime}_{ij}$ form a basis of $H^{0}(\mathcal{G}_{ij})$. We set
$V=\operatorname{Image}{(H^{0}(\Omega_{C}^{1}\otimes\eta\otimes\mathcal{O}_{C}(-E_{\infty}))\hookrightarrow{H^{0}(\Omega_{C}^{1}\otimes\eta)})}$
and
$V_{k}=\operatorname{Image}{(H^{0}(\Omega_{C}^{1}\otimes\eta\otimes\mathcal{O}_{C}(-E_{\infty}-p_{k}))\hookrightarrow H^{0}(\Omega_{C}^{1}\otimes\eta))}$
for $1\leq{k}\leq{g+1}$. Since $\dim{V}=g+1$, by (1) and (5), we have $\bigcap_{k=1}^{g+1}V_{k}=\{0\}$, and $\dim{\bigl(\bigcap_{k\neq{i}}V_{k}\bigr)}=1$ for $1\leq{i}\leq{g+1}$. Thus, there is a unique divisor $D_{i}\in|\Omega_{C}^{1}\otimes\eta\otimes\mathcal{H}^{\vee}\otimes\mathcal{O}_{C}(-p_{1}-\dots-p_{g+1}+p_{i})|$. By (1) and (5), we have
$D_{i}+p_{j}\neq{D_{j}+p_{i}}\in|\Omega_{C}^{1}\otimes\eta\otimes\mathcal{G}_{ij}^{\vee}|.$
for $1\leq{i}<j\leq{g+1}$. Let $\beta_{ij}\in{H^{0}(\Omega_{C}^{1}\otimes\eta\otimes\mathcal{G}_{ij}^{\vee})}$ define $D_{i}+p_{j}$, and $\beta^{\prime}_{ij}\in{H^{0}(\Omega_{C}^{1}\otimes\eta\otimes\mathcal{G}_{ij}^{\vee})}$ define $D_{j}+p_{i}$. By (3), $\beta_{ij},\beta^{\prime}_{ij}$ form a basis of $H^{0}(\Omega_{C}^{1}\otimes\eta\otimes\mathcal{G}_{ij}^{\vee})$. Let $\gamma_{i}$ be a generator of $\bigcap_{k\neq{i}}V_{k}$ for $1\leq{i}\leq{g+1}$. Then $\gamma_{1},\dots,\gamma_{g+1}$ form a basis of $V$. Since $\mu(\alpha_{ij},\beta_{ij})\in\bigcap_{k\neq{i}}V_{k}$ and $\mu(\alpha_{ij},\beta^{\prime}_{ij})\in\bigcap_{k\neq{j}}V_{k}$, we have $\mu(\alpha_{ij},\beta_{ij})=c_{ij}\gamma_{i}$ and $\mu(\alpha_{ij},\beta^{\prime}_{ij})=c^{\prime}_{ij}\gamma_{j}$ for some $c_{ij},c^{\prime}_{ij}\in\mathbb{C}^{*}$. By using the rational function $f$, we define the injective homomorphism
$\lambda:\,V\longrightarrow{H^{0}(\Omega_{C}^{1}\otimes\eta)};\,\gamma\longmapsto{f\gamma}.$
Then, by [1, p195, Corollary 2], the defining quadratics
$\det{\begin{pmatrix}\mu(\alpha_{ij},\beta_{ij})&\mu(\alpha_{ij},\beta^{\prime}_{ij})\\ \mu(\alpha^{\prime}_{ij},\beta_{ij})&\mu(\alpha^{\prime}_{ij},\beta^{\prime}_{ij})\\ \end{pmatrix}}=\det{\begin{pmatrix}\mu(\alpha_{ij},\beta_{ij})&\mu(\alpha_{ij},\beta^{\prime}_{ij})\\ \lambda(\mu(\alpha_{ij},\beta_{ij}))&\lambda(\mu(\alpha_{ij},\beta^{\prime}_{ij}))\\ \end{pmatrix}}$
$=$ $\det{\begin{pmatrix}c_{ij}\gamma_{i}&c^{\prime}_{ij}\gamma_{j}\\ \lambda(c_{ij}\gamma_{i})&\lambda(c^{\prime}_{ij}\gamma_{j})\\ \end{pmatrix}}=c_{ij}c^{\prime}_{ij}\det{\begin{pmatrix}\gamma_{i}&\gamma_{j}\\ \lambda(\gamma_{i})&\lambda(\gamma_{j})\\ \end{pmatrix}}\in\operatorname{Sym}^{2}{H^{0}(\Omega_{C}^{1}\otimes\eta)}$
of $Q_{0,\mathcal{G}_{ij}}$ for $1\leq{i}<{j}\leq{g+1}$ are linearly independent over $\mathbb{C}$. Since
$\dim{\operatorname{Ker}{(\rho_{2})}}=\frac{(g+3)(g+4)}{2}-(3g+5)=\frac{g(g+1)}{2}+1,$
the defining quadratics of $Q_{0,\mathcal{G}_{ij}}$ for $1\leq{i}<j\leq{g+1}$ generate a subspace of codimension $1$ in $\operatorname{Ker}{(\rho_{2})}$.
The quadric hypersurface $Q_{0,\mathcal{G}_{ij}}\subset\mathbb{P}{(H^{0}(\Omega_{C}^{1}\otimes\eta))}$ contains the line
$\ell_{V}=\mathbb{P}{(H^{0}(\Omega_{C}^{1}\otimes\eta)/V)}\subset\mathbb{P}{(H^{0}(\Omega_{C}^{1}\otimes\eta))}$
for $1\leq{i}<j\leq{g+1}$. It is enough to show that $\ell_{V}$ is not contained in
$Q_{0,\mathcal{G}_{0}}=\bigcup_{D\in|\mathcal{G}_{0}|}\mathbb{P}(H^{0}(\Omega_{C}^{1}\otimes\eta)/V_{D})\subset\mathbb{P}{(H^{0}(\Omega_{C}^{1}\otimes\eta))}.$
If $D\in|\mathcal{G}_{0}|$ satisfies $\operatorname{supp}{(E_{\infty})}\cap\operatorname{supp}{(D)}=\emptyset$, then, by (5), we have $V_{D}\cap{V}=\{0\}$ and $\ell_{V}\cap\mathbb{P}(H^{0}(\Omega_{C}^{1}\otimes\eta)/V_{D})=\emptyset$. By (2) and (4), since the pencil $|\mathcal{G}_{0}|$ has no base points, there exist at most two $D\in|\mathcal{G}_{0}|$ such that $\operatorname{supp}{(E_{\infty})}\cap\operatorname{supp}{(D)}\neq\emptyset$. When $p\in\operatorname{supp}{(E_{\infty})}\cap\operatorname{supp}{(D)}$ and $E_{\infty}=p+p^{\prime}$, by (4), we have $p^{\prime}\nleq{D}$. If $V_{D}\subset{V}$, then
$2=h^{0}(\Omega_{C}^{1}\otimes\eta\otimes\mathcal{O}_{C}(-D))=h^{0}(\Omega_{C}^{1}\otimes\eta\otimes\mathcal{O}_{C}(-D-p^{\prime})),$
and $h^{0}(\Omega_{C}^{1}\otimes\eta\otimes\mathcal{O}_{C}(-D-p-p^{\prime}))\geq 1$. But this contradicts (5), hence $V_{D}\nsubseteq{V}$ and $\ell_{V}\cap\mathbb{P}(H^{0}(\Omega_{C}^{1}\otimes\eta)/V_{D})=\{\Phi(p)\}$. This implies that $\ell_{V}\nsubseteq{Q_{0,\mathcal{G}_{0}}}$. ∎
Proof of Proposition 3.1.
We prove it by induction on $n$. When $n=4$, the result holds by Lemma 3.2. We assume $n\geq 5$. Let $p_{1},p_{2},p_{3}\in{C\setminus\Sigma}$ be three distinct points. We denote by
$\iota_{i}:\,\operatorname{Sym}^{2}{H^{0}(\Omega_{C}^{1}\otimes\eta\otimes\mathcal{O}_{C}(-p_{i}))}\hookrightarrow\operatorname{Sym}^{2}{H^{0}(\Omega_{C}^{1}\otimes\eta)}$
the natural injective homomorphism. Since the multiplication maps
$\rho_{2,p_{i}}:\,\operatorname{Sym}^{2}{H^{0}(\Omega_{C}^{1}\otimes\eta\otimes\mathcal{O}_{C}(-p_{i}))}\longrightarrow H^{0}((\Omega_{C}^{1}\otimes\eta\otimes\mathcal{O}_{C}(-p_{i}))^{\otimes 2})$
and
$\rho_{2,p_{i}+p_{j}}:\,\operatorname{Sym}^{2}{H^{0}(\Omega_{C}^{1}\otimes\eta\otimes\mathcal{O}_{C}(-p_{i}-p_{j}))}\longrightarrow H^{0}((\Omega_{C}^{1}\otimes\eta\otimes\mathcal{O}_{C}(-p_{i}-p_{j}))^{\otimes 2})$
are surjective, we have
$\dim{(\iota_{1}(\operatorname{Ker}{(\rho_{2,p_{1}})})+\iota_{2}(\operatorname{Ker}{(\rho_{2,p_{2}})}))}$
$=$ $\dim{\iota_{1}(\operatorname{Ker}{(\rho_{2,p_{1}})})}+\dim{\iota_{2}(\operatorname{Ker}{(\rho_{2,p_{2}})})}-\dim{(\iota_{1}(\operatorname{Ker}{(\rho_{2,p_{1}})})\cap\iota_{2}(\operatorname{Ker}{(\rho_{2,p_{2}})}))}$
$=$ $\dim{\operatorname{Ker}{(\rho_{2,p_{1}})}}+\dim{\operatorname{Ker}{(\rho_{2,p_{2}})}}-\dim{\operatorname{Ker}{(\rho_{2,p_{1}+p_{2}})}}=\dim{\operatorname{Ker}{(\rho_{2})}}-1.$
Let
$\pi_{i}:\,\mathbb{P}(H^{0}(\Omega_{C}^{1}\otimes\eta))\setminus\{\Phi(p_{i})\}\longrightarrow\mathbb{P}(H^{0}(\Omega_{C}^{1}\otimes\eta\otimes\mathcal{O}_{C}(-p_{i})))$
denote the projection from $\Phi(p_{i})$, and let
$\Phi_{i}=\Phi_{|\Omega_{C}^{1}\otimes\eta\otimes\mathcal{O}_{C}(-p_{i})|}:\,C\longrightarrow\mathbb{P}(H^{0}(\Omega_{C}^{1}\otimes\eta\otimes\mathcal{O}_{C}(-p_{i})))$
be the closed immersion defined by the very ample linear system $|\Omega_{C}^{1}\otimes\eta\otimes\mathcal{O}_{C}(-p_{i})|$. Let $\ell\subset\mathbb{P}(H^{0}(\Omega_{C}^{1}\otimes\eta))$ be the line through the points $\Phi(p_{1})$ and $\Phi(p_{2})$. Since
$\pi_{1}(\ell)=\pi_{1}(\Phi(p_{2}))=\Phi_{1}(p_{2})\neq\Phi_{1}(p_{3})=\pi_{1}(\Phi(p_{3})),$
we have $\Phi(p_{3})\notin\ell$. Hence, $\pi_{3}(\ell)$ is a line in $\mathbb{P}(H^{0}(\Omega_{C}^{1}\otimes\eta\otimes\mathcal{O}_{C}(-p_{3})))$, and we have
$\pi_{3}(\ell)\nsubseteq\Phi_{3}(C)\subset\mathbb{P}(H^{0}(\Omega_{C}^{1}\otimes\eta\otimes\mathcal{O}_{C}(-p_{3}))).$
By [15], we have $\Phi_{3}(C)=\bigcap_{g\in\operatorname{Ker}{(\rho_{2,p_{3}})}}Q_{g}$, where $Q_{g}\subset\mathbb{P}(H^{0}(\Omega_{C}^{1}\otimes\eta\otimes\mathcal{O}_{C}(-p_{3})))$ denotes the quadric defined by $g$. Thus, there exists $g\in\operatorname{Ker}{(\rho_{2,p_{3}})}$ such that $\pi_{3}(\ell)\nsubseteq{Q_{g}}$. Since the quadric $\overline{\pi_{3}^{-1}(Q_{g})}\subset\mathbb{P}(H^{0}(\Omega_{C}^{1}\otimes\eta))$ defined by $\iota_{3}(g)\in\operatorname{Ker}{(\rho_{2})}$ does not contain the line $\ell$, we have $\iota_{3}(g)\notin\iota_{1}(\operatorname{Ker}{(\rho_{2,p_{1}})})+\iota_{2}(\operatorname{Ker}{(\rho_{2,p_{2}})})$. Hence,
$\operatorname{Ker}{(\rho_{2})}=\iota_{1}(\operatorname{Ker}{(\rho_{2,p_{1}})})+\iota_{2}(\operatorname{Ker}{(\rho_{2,p_{2}})})+\iota_{3}(\operatorname{Ker}{(\rho_{2,p_{3}})}).$
By the induction assumption, $\operatorname{Ker}{(\rho_{2,p_{i}})}\subset\operatorname{Sym}^{2}{H^{0}(\Omega_{C}^{1}\otimes\eta\otimes\mathcal{O}_{C}(-p_{i}))}$ is generated by the defining quadratics of $Q_{E,\mathcal{F}}\subset\mathbb{P}(H^{0}(\Omega_{C}^{1}\otimes\eta\otimes\mathcal{O}_{C}(-p_{i})))$ for $E\in{(C\setminus\Sigma)^{(n-5)}}$ and $\mathcal{F}\in{K_{2}(E,\eta\otimes\mathcal{O}_{C}(-p_{i}))}$. Hence, $\iota_{i}(\operatorname{Ker}{(\rho_{2,p_{i}})})$ is generated by the defining quadratics of $Q_{E+p_{i},\mathcal{F}}=\overline{\pi_{i}^{-1}(Q_{E,\mathcal{F}})}\subset\mathbb{P}(H^{0}(\Omega_{C}^{1}\otimes\eta))$ for $E+p_{i}\in{(C\setminus\Sigma)^{(n-4)}}$ and $\mathcal{F}\in{K_{2}(E,\eta\otimes\mathcal{O}_{C}(-p_{i}))=K_{2}(E+p_{i},\eta)}$. ∎
Corollary 3.3.
Let $(P,[\mathcal{L}])$ be the Prym variety of $\tilde{C}\overset{\phi}{\rightarrow}{C}]\in\mathcal{R}_{g,2n}$. If $n\geq 4$ and $C$ is an elliptic curve or a hyperelliptic curve, then $\operatorname{Ker}{(\rho_{2})}$ is generated by the defining quadratics of the tangent cones $\mathcal{T}_{\Theta,x}$ for $x\in{S(P,\mathcal{L})}$ and $\Theta\in|\mathcal{L}|$ satisfying $\operatorname{mult}_{x}{\Theta}=2$.
Proof.
Since $S(P,\mathcal{L})\overset{\simeq}{\rightarrow}S_{2}(P_{\phi},\mathcal{L}_{\phi})$, by Proposition 3.1, it is enough to show that for $E\in{(C\setminus\operatorname{Branch}{(\phi)})^{(n-4)}}$ and $\mathcal{G}\in{K_{2}(E,\eta_{\phi})}$, the quadric $\mathcal{V}_{E,\mathcal{G}}\subset{H^{0}(\Omega_{C}^{1}\otimes\eta_{\phi})^{\vee}}$ is the tangent cone $\mathcal{T}_{W_{\xi},\mathcal{F}}$ for some $\mathcal{F}\in{S_{2}(P_{\phi},\mathcal{L}_{\phi})}$ and $\xi\in\operatorname{Pic}^{0}{(C)}$ satisfying $\operatorname{mult}_{\mathcal{F}}{W_{\xi}}=2$. For $E\in{(C\setminus\operatorname{Branch}{(\phi)})^{(n-4)}}$, there exist $D\in\tilde{C}^{(n-4)}$ such that $\phi(D)=E$ and $\phi^{*}p\nleq{D}$ for any $p\in{C}$. Let $\delta\in\operatorname{Pic}^{g+1}{(C)}$ be an invertible sheaf such that $\delta^{\otimes 2}=\Omega_{C}^{1}\otimes\eta_{\phi}\otimes\mathcal{O}_{C}(-E)$. For $\mathcal{G}\in{K_{2}(E,\eta_{\phi})}$, we set $\mathcal{F}=\mathcal{O}_{\tilde{C}}(D)\otimes\phi^{*}\delta$ and $\xi=\mathcal{G}\otimes\delta^{\vee}\in\operatorname{Pic}^{0}{(C)}$. Then, by Lemma 2.6, we have $\operatorname{mult}_{\mathcal{F}}{W_{\xi}}=2$, and $\mathcal{V}_{E,\mathcal{G}}$ is the tangent cone of $W_{\xi}$ at $\mathcal{F}$. ∎
4. Andreotti-Mayer loci
Let
$\mathfrak{H}_{d}=\{\tau\in\operatorname{Mat}_{d}{(\mathbb{C})}\mid\tau=\operatorname{\mathstrut{}^{\mathit{t}}\!}{\tau},\,\operatorname{Im}{(\tau)}>0\}$
be the Siegel upper half-space of degree d=g+n1d=g+n-1. For the type of the polarization $\Delta=(\underbrace{1,\dots,1}_{n-1},\underbrace{2,\dots,2}_{g})$, we denote the diagonal matrix $\operatorname{diag}{\Delta}\in\operatorname{Mat}_{d}{(\mathbb{Z})}$ by the same notation $\Delta$. We define the action of $\mathbb{Z}^{d}\times\mathbb{Z}^{d}$ on $\mathbb{C}^{d}\times\mathfrak{H}_{d}$ by
$(\mathbb{Z}^{d}\times\mathbb{Z}^{d})\times(\mathbb{C}^{d}\times\mathfrak{H}_{d})\longrightarrow\mathbb{C}^{d}\times\mathfrak{H}_{d};\,(({\boldsymbol{m}}^{\prime},{\boldsymbol{m}}^{\prime\prime}),({\boldsymbol{z}},\tau))\longmapsto({\boldsymbol{z}}+\tau{\boldsymbol{m}}^{\prime}+\Delta{\boldsymbol{m}}^{\prime\prime},\tau),$
and denote its quotient by $\mathcal{U}_{d}^{\Delta}$. Then
$u:\,\mathcal{U}_{d}^{\Delta}\longrightarrow\mathfrak{H}_{d};\,({\boldsymbol{z}},\tau)\longmapsto\tau$
is a proper morphism of complex manifolds, and the fiber $A_{\tau}=u^{-1}(\tau)$ at $\tau\in\mathfrak{H}_{d}$ is an abelian variety of dimension $d$. Let $\mathcal{L}_{\tau}$ be the invertible sheaf on $A_{\tau}$ defined by the Hermitian form
$H:\,\mathbb{C}^{d}\times\mathbb{C}^{d}\longrightarrow\mathbb{C};\,({\boldsymbol{z}},{\boldsymbol{w}})\longmapsto\operatorname{\mathstrut{}^{\mathit{t}}\!}{{\boldsymbol{z}}}\,(\operatorname{Im}{(\tau)})^{-1}\,\overline{\boldsymbol{w}}$
and the semi-character
$\chi_{0}:\,\tau\mathbb{Z}^{d}+\Delta\mathbb{Z}^{d}\longrightarrow\mathbb{C}^{*};\,\tau{\boldsymbol{m}}^{\prime}+\Delta{\boldsymbol{m}}^{\prime\prime}\longmapsto\exp{\bigl(\pi\sqrt{-1}\,\operatorname{\mathstrut{}^{\mathit{t}}\!}{\boldsymbol{m}}^{\prime}\Delta{\boldsymbol{m}}^{\prime\prime}\bigr)}.$
Then $(A_{\tau},[\mathcal{L}_{\tau}])$ is a polarized abelian variety of type $\Delta$. The moduli space $\mathcal{A}_{d}^{\Delta}$ is the quotient of $\mathfrak{H}_{d}$ by the action
$\Gamma_{\Delta}\times\mathfrak{H}_{d}\longrightarrow\mathfrak{H}_{d};\,(\begin{pmatrix}\alpha&\beta\\ \gamma&\delta\end{pmatrix},\tau)\longmapsto(\alpha\tau+\beta)(\gamma\tau+\delta)^{-1},$
where $\Gamma_{\Delta}$ is the discrete subgroup of $\operatorname{Sp}_{2d}{(\mathbb{R})}$ defined by
$\Gamma_{\Delta}=\left\{\begin{pmatrix}\alpha&\beta\\ \gamma&\delta\end{pmatrix}\in\operatorname{Sp}_{2d}{(\mathbb{R})}\,\middle|\,\begin{pmatrix}\alpha&\beta{\Delta^{-1}}\\ \Delta\gamma&\Delta\delta{\Delta^{-1}}\end{pmatrix}\in\operatorname{Mat}_{2d}{(\mathbb{Z})}\right\}.$
Following [12], for ${\boldsymbol{c}}^{\prime},{\boldsymbol{c}}^{\prime\prime}\in\frac{1}{2}\mathbb{Z}^{d}$ the theta function is defined by
$\theta\begin{bmatrix}\operatorname{\mathstrut{}^{\mathit{t}}\!}{\boldsymbol{c}}^{\prime}\\ \operatorname{\mathstrut{}^{\mathit{t}}\!}{\boldsymbol{c}}^{\prime\prime}\end{bmatrix}({\boldsymbol{z}},\tau)=\sum_{{\boldsymbol{m}}\in\mathbb{Z}^{d}}\exp{\Bigl(\pi\sqrt{-1}\bigl(\operatorname{\mathstrut{}^{\mathit{t}}\!}{({\boldsymbol{m}}+{\boldsymbol{c}}^{\prime})}\tau({\boldsymbol{m}}+{\boldsymbol{c}}^{\prime})+2\operatorname{\mathstrut{}^{\mathit{t}}\!}{({\boldsymbol{m}}+{\boldsymbol{c}}^{\prime})}({\boldsymbol{z}}+{\boldsymbol{c}}^{\prime\prime})\bigr)\Bigr)}.$
Then, by [2, Theorem 3.2.7. and Remark 8.5.3.],
$\theta_{\boldsymbol{i}}({\boldsymbol{z}},\tau)=\theta\begin{bmatrix}0&\cdots&0&i_{1}&\cdots&i_{g}\\ 0&\cdots&0&0&\cdots&0\end{bmatrix}({\boldsymbol{z}},\tau),\quad\Bigl({\boldsymbol{i}}=\begin{pmatrix}i_{1}\\ \vdots\\ i_{g}\end{pmatrix}\in{I=\Bigl\{0,\frac{1}{2}\Bigr\}^{g}}\Bigr)$
form a basis of $H^{0}(A_{\tau},\mathcal{L}_{\tau})$. We consider the complex analytic space
$\mathcal{S}=\bigl\{({\boldsymbol{z}},\tau)\in\mathcal{U}_{d}^{\Delta}\mid\theta_{\boldsymbol{i}}({\boldsymbol{z}},\tau)=0,\,\frac{\partial\theta_{\boldsymbol{i}}}{\partial{z_{j}}}({\boldsymbol{z}},\tau)=0\quad({\boldsymbol{i}}\in{I},\,1\leq{j}\leq{d})\bigr\},$
where $z_{j}$ denotes the coordinate of ${\boldsymbol{z}}=\begin{pmatrix}z_{1}\\ \vdots\\ z_{d}\end{pmatrix}\in\mathbb{C}^{d}$. Let $\upsilon=u|_{\mathcal{S}}:\mathcal{S}\rightarrow\mathfrak{H}_{d}$ be the restriction of $u:\mathcal{U}_{d}^{\Delta}\rightarrow\mathfrak{H}_{d}$. Then the fiber of $\upsilon$ at $\tau\in\mathfrak{H}_{d}$ is
$\upsilon^{-1}(\tau)=S(A_{\tau},\mathcal{L}_{\tau})=\bigcap_{\Theta\in|\mathcal{L}_{\tau}|}\Theta_{\mathrm{sing}}.$
Lemma 4.1.
If $({\boldsymbol{a}},\tau)\in\mathcal{S}$, then $\bigl({\boldsymbol{a}}+\tau\begin{pmatrix}{\boldsymbol{0}}\\ {\boldsymbol{i}}\end{pmatrix}+\Delta\begin{pmatrix}{\boldsymbol{0}}\\ {\boldsymbol{j}}\end{pmatrix},\tau\bigr)\in\mathcal{S}$ for ${\boldsymbol{i}},{\boldsymbol{j}}\in{I}$.
Proof.
For ${\boldsymbol{i}},{\boldsymbol{i}}^{\prime}\in{I}$, there exists ${\boldsymbol{i}}^{\prime\prime}\in{I}$ such that ${\boldsymbol{i}}+{\boldsymbol{i}}^{\prime}-{\boldsymbol{i}}^{\prime\prime}\in\mathbb{Z}^{g}$. We set ${\boldsymbol{c}}=\begin{pmatrix}c_{1}\\ \vdots\\ c_{d}\end{pmatrix}=\begin{pmatrix}{\boldsymbol{0}}\\ {\boldsymbol{i}}\end{pmatrix}\in\frac{1}{2}\mathbb{Z}^{d}$ and ${\boldsymbol{d}}=\Delta\begin{pmatrix}{\boldsymbol{0}}\\ {\boldsymbol{j}}\end{pmatrix}\in\mathbb{Z}^{d}$. Then we have
$\theta_{{\boldsymbol{i}}^{\prime}}({\boldsymbol{z}}+\tau{\boldsymbol{c}}+{\boldsymbol{d}},\tau)=\exp{\Bigl(\pi\sqrt{-1}(-\operatorname{\mathstrut{}^{\mathit{t}}\!}{\boldsymbol{c}}\tau{\boldsymbol{c}}-2\operatorname{\mathstrut{}^{\mathit{t}}\!}{\boldsymbol{c}}{\boldsymbol{z}}+4\operatorname{\mathstrut{}^{\mathit{t}}\!}{\boldsymbol{i}}^{\prime}{\boldsymbol{j}})\Bigr)}\theta_{{\boldsymbol{i}}^{\prime\prime}}({\boldsymbol{z}},\tau),$
and
$\frac{\partial{\theta_{{\boldsymbol{i}}^{\prime}}}}{\partial{z_{j}}}({\boldsymbol{z}}+\tau{\boldsymbol{c}}+{\boldsymbol{d}},\tau)\\ =\exp{\Bigl(\pi\sqrt{-1}(-\operatorname{\mathstrut{}^{\mathit{t}}\!}{\boldsymbol{c}}\tau{\boldsymbol{c}}-2\operatorname{\mathstrut{}^{\mathit{t}}\!}{\boldsymbol{c}}{\boldsymbol{z}}+4\operatorname{\mathstrut{}^{\mathit{t}}\!}{\boldsymbol{i}}^{\prime}{\boldsymbol{j}})\Bigr)}\bigl(-2\pi\sqrt{-1}\,c_{j}\theta_{{\boldsymbol{i}}^{\prime\prime}}({\boldsymbol{z}},\tau)+\frac{\partial{\theta_{{\boldsymbol{i}}^{\prime\prime}}}}{\partial{z_{j}}}({\boldsymbol{z}},\tau)\bigr)$
for $1\leq{j}\leq{d}$. Since $\theta_{{\boldsymbol{i}}^{\prime\prime}}({\boldsymbol{a}},\tau)=0$ and $\frac{\partial{\theta_{{\boldsymbol{i}}^{\prime\prime}}}}{\partial{z_{j}}}({\boldsymbol{a}},\tau)=0$, we have $\theta_{{\boldsymbol{i}}^{\prime}}({\boldsymbol{a}}+\tau{\boldsymbol{c}}+{\boldsymbol{d}},\tau)=0$ and $\frac{\partial{\theta_{{\boldsymbol{i}}^{\prime}}}}{\partial{z_{j}}}({\boldsymbol{a}}+\tau{\boldsymbol{c}}+{\boldsymbol{d}},\tau)=0$. ∎
For $m\geq 0$, the degeneracy set ([5, 3.6.])
$\mathcal{S}_{m}=\{({\boldsymbol{z}},\tau)\in\mathcal{S}\mid\dim_{({\boldsymbol{z}},\tau)}{\upsilon^{-1}(\tau)}\geq{m}\}$
is an analytic subset in $\mathcal{S}$, and by the proper mapping theorem ([6, 10.6.1.]), $\mathcal{N}_{m}=\upsilon(\mathcal{S}_{m})$ is an analytic subset in $\mathfrak{H}_{d}$. Let $\mathcal{N}$ be a local irreducible component of $\mathcal{N}_{m}$ at $\tau_{0}\in\mathcal{N}_{m}$. We denote by $\Lambda(\mathcal{N})$ the set of all irreducible components $\mathcal{M}$ of $\upsilon^{-1}(\mathcal{N})\cap\mathcal{S}_{m}$ such that $\upsilon(\mathcal{M})=\mathcal{N}$. Since $\upsilon$ is proper, $\Lambda(\mathcal{N})$ is a non-empty finite set. Following [4], we consider the subset $\mathcal{S}(\tau_{0},\mathcal{N})=\upsilon^{-1}(\tau_{0})\cap\bigcup_{\mathcal{M}\in\Lambda(\mathcal{N})}\mathcal{M}$ in $\upsilon^{-1}(\tau_{0})$.
Lemma 4.2.
If $\mathcal{N}$ is nonsingular, then
$T_{\mathcal{N},\tau_{0}}\subset\Bigl(\sum_{1\leq{k}\leq{l}\leq{d}}\frac{\partial\theta_{\boldsymbol{i}}}{\partial\tau_{kl}}({\boldsymbol{a}},\tau_{0})\,d\tau_{kl}\Bigr)^{\perp}\subset{T_{\mathfrak{H}_{d},\tau_{0}}}$
for ${\boldsymbol{i}}\in{I}$ and $({\boldsymbol{a}},\tau_{0})\in{\mathcal{S}(\tau_{0},\mathcal{N})}$, where $\tau_{kl}=\tau_{lk}$ denotes the coordinate of $\tau=\begin{pmatrix}\tau_{11}&&\tau_{1d}\\ &\cdots&\\ \tau_{d1}&&\tau_{dd}\end{pmatrix}\in\mathfrak{H}_{d}$.
Proof.
Since $({\boldsymbol{a}},\tau_{0})\in{\mathcal{S}(\tau_{0},\mathcal{N})}$, there exists $\mathcal{M}\in\Lambda(\mathcal{N})$ such that $({\boldsymbol{a}},\tau_{0})\in\mathcal{M}$. Then,
$U=\{({\boldsymbol{z}},\tau)\in\mathcal{M}\setminus\mathcal{M}_{\mathrm{sing}}\mid\text{$T_{\mathcal{M},({\boldsymbol{z}},\tau)}\overset{d\upsilon}{\rightarrow}T_{\mathcal{N},\tau}$ is surjective}\}$
is a non-empty open subset of $\mathcal{M}$. For $({\boldsymbol{z}},\tau)\in{U}$, we can take a local coordinate ${\boldsymbol{t}}=(t_{1},\dots,t_{\dim{\mathcal{N}}})$ of $\mathcal{N}$ at $\tau$ and a local coordinate $({\boldsymbol{s}},{\boldsymbol{t}})$ of $U$ at $({\boldsymbol{z}},\tau)$ such that $\upsilon|_{U}:U\rightarrow\mathcal{N}$ is given by $({\boldsymbol{s}},{\boldsymbol{t}})\mapsto{\boldsymbol{t}}$. Let $({\boldsymbol{z}}({\boldsymbol{s}},{\boldsymbol{t}}),\tau({\boldsymbol{t}}))\in{U}\subset\mathcal{S}$ be the point corresponding to $({\boldsymbol{s}},{\boldsymbol{t}})$. Then we have
$\theta_{\boldsymbol{i}}({\boldsymbol{z}}({\boldsymbol{s}},{\boldsymbol{t}}),\tau({\boldsymbol{t}}))=0,\quad\frac{\partial\theta_{\boldsymbol{i}}}{\partial{z_{j}}}({\boldsymbol{z}}({\boldsymbol{s}},{\boldsymbol{t}}),\tau({\boldsymbol{t}}))=0\quad({\boldsymbol{i}}\in{I},\,1\leq{j}\leq{d}).$
For $1\leq{\nu}\leq\dim{\mathcal{N}}$, by the chain rule, we get
$\frac{\partial}{\partial{t_{\nu}}}\Bigl(\theta_{\boldsymbol{i}}({\boldsymbol{z}}({\boldsymbol{s}},{\boldsymbol{t}}),\tau({\boldsymbol{t}}))\Bigr)\\ =\sum_{j=1}^{d}\frac{\partial\theta_{\boldsymbol{i}}}{\partial{z_{j}}}({\boldsymbol{z}}({\boldsymbol{s}},{\boldsymbol{t}}),\tau({\boldsymbol{t}}))\frac{\partial{z_{j}}}{\partial{t_{\nu}}}({\boldsymbol{s}},{\boldsymbol{t}})+\sum_{1\leq{k}\leq{l}\leq{d}}\frac{\partial\theta_{\boldsymbol{i}}}{\partial{\tau_{kl}}}({\boldsymbol{z}}({\boldsymbol{s}},{\boldsymbol{t}}),\tau({\boldsymbol{t}}))\frac{\partial{\tau_{kl}}}{\partial{t_{\nu}}}({\boldsymbol{t}}),$
which implies
$\sum_{1\leq{k}\leq{l}\leq{d}}\frac{\partial\theta_{\boldsymbol{i}}}{\partial{\tau_{kl}}}({\boldsymbol{z}}({\boldsymbol{s}},{\boldsymbol{t}}),\tau({\boldsymbol{t}}))\frac{\partial{\tau_{kl}}}{\partial{t_{\nu}}}({\boldsymbol{t}})=0.$
This shows that
$\frac{\partial}{\partial{t_{\nu}}}=\sum_{1\leq{k}\leq{l}\leq{d}}\frac{\partial{\tau_{kl}}}{\partial{t_{\nu}}}({\boldsymbol{t}})\frac{\partial}{\partial{\tau_{kl}}}\in\Bigl(\sum_{1\leq{k}\leq{l}\leq{d}}\frac{\partial\theta_{\boldsymbol{i}}}{\partial{\tau_{kl}}}({\boldsymbol{z}}({\boldsymbol{s}},{\boldsymbol{t}}),\tau({\boldsymbol{t}}))\,d\tau_{kl}\Bigr)^{\perp}\subset{T_{\mathfrak{H}_{d},\tau({\boldsymbol{t}})}}.$
Since the open subset $U\subset\mathcal{M}$ is contained in the analytic subset
$\mathcal{M}^{\prime}=\left\{({\boldsymbol{z}},\tau)\in\mathcal{M}\,\middle|\,T_{\mathcal{N},\tau}\subset\Bigl(\sum_{1\leq{k}\leq{l}\leq{d}}\frac{\partial\theta_{\boldsymbol{i}}}{\partial\tau_{kl}}({\boldsymbol{z}},\tau)\,d\tau_{kl}\Bigr)^{\perp}\right\},$
we have $({\boldsymbol{a}},\tau_{0})\in\mathcal{M}=\mathcal{M}^{\prime}$. ∎
Proposition 4.3.
Assume that $\dim{\upsilon^{-1}(\tau_{0})}=m$. If there exist $c_{{\boldsymbol{i}},j}\in\mathbb{C}$ and $({\boldsymbol{a}}_{j},\tau_{0})\in\mathcal{S}(\tau_{0},\mathcal{N})$ for ${\boldsymbol{i}}\in{I}$ and $1\leq{j}\leq{r}$ such that cotangent vectors
$\sum_{1\leq{k}\leq{l}\leq{d}}\sum_{{\boldsymbol{i}}\in{I}}c_{{\boldsymbol{i}},j}\frac{\partial\theta_{\boldsymbol{i}}}{\partial\tau_{kl}}({\boldsymbol{a}}_{j},\tau_{0})\,d\tau_{kl}\in{T_{\mathfrak{H}_{d},\tau_{0}}^{\vee}}=\bigoplus_{1\leq{k}\leq{l}\leq{d}}\mathbb{C}d\tau_{kl}\quad(1\leq{j}\leq{r})$
are linearly independent, then $\dim{\mathcal{N}}\leq\dim{\mathfrak{H}_{d}}-r$.
Proof.
For $1\leq{j}\leq{r}$, let $\mathcal{M}_{j}\in\Lambda(\mathcal{N})$ satisfy $({\boldsymbol{a}}_{j},\tau_{0})\in\mathcal{M}_{j}$. By the upper semi-continuity of the fiber dimension [5, 3.4.], there exist open neighborhoods $U_{j}\subset\mathcal{M}_{j}$ at $({\boldsymbol{a}}_{j},\tau_{0})$ such that
$m=\dim{\upsilon^{-1}(\tau_{0})}\geq\dim_{({\boldsymbol{a}}_{j},\tau_{0})}{(\upsilon^{-1}(\tau_{0})\cap\mathcal{M})}\geq\dim_{({\boldsymbol{b}},\tau)}{(\upsilon^{-1}(\tau)\cap\mathcal{M})}\geq{m}$
for any $({\boldsymbol{b}},\tau)\in{U_{j}}$. Then, by [5, 3.10.], the restriction $\upsilon|_{U_{j}}:U_{j}\rightarrow\mathcal{N}$ is an open mapping. By replacing $U_{j}$ with smaller open neighborhoods, we may assume that
$\sum_{1\leq{k}\leq{l}\leq{d}}\sum_{{\boldsymbol{i}}\in{I}}c_{{\boldsymbol{i}},j}\frac{\partial\theta_{\boldsymbol{i}}}{\partial\tau_{kl}}({\boldsymbol{b}}_{j},\tau_{j})\,d\tau_{kl}\quad(1\leq{j}\leq{r})$
are linearly independent for any $\{({\boldsymbol{b}}_{j},\tau_{j})\}_{1\leq{j}\leq{r}}\in\prod_{j=1}^{r}{U_{j}}$. Since $\bigcap_{j=1}^{r}\upsilon(U_{j})$ is an open neighborhood of $\mathcal{N}$ at $\tau_{0}$, there exists a point $\tau^{\prime}_{0}\in\bigcap_{j=1}^{r}\upsilon(U_{j})$ such that $\mathcal{N}$ is nonsingular at $\tau^{\prime}_{0}$. Then, $\mathcal{N}^{\prime}=\mathcal{N}\setminus\mathcal{N}_{\mathrm{sing}}$ is a local irreducible component of $\mathcal{N}_{m}$ at $\tau_{0}^{\prime}$. For $({\boldsymbol{a}}^{\prime}_{j},\tau^{\prime}_{0})\in\upsilon^{-1}(\tau^{\prime}_{0})\cap{U_{j}}$, we have $({\boldsymbol{a}}^{\prime}_{j},\tau^{\prime}_{0})\in\mathcal{S}(\tau^{\prime}_{0},\mathcal{N}^{\prime})$ and
$\sum_{1\leq{k}\leq{l}\leq{d}}\sum_{{\boldsymbol{i}}\in{I}}c_{{\boldsymbol{i}},j}\frac{\partial\theta_{\boldsymbol{i}}}{\partial\tau_{kl}}({\boldsymbol{a}}^{\prime}_{j},\tau^{\prime}_{0})\,d\tau_{kl}\quad(1\leq{j}\leq{r})$
are linearly independent. By Lemma 4.2, we conclude that $\dim{\mathcal{N}}\leq\dim{\mathfrak{H}_{d}}-r$. ∎
[…]
Lemma 4.4.
If $n\geq 4$ and $\tau_{0}\in\pi^{-1}(\operatorname{Prym}_{g,2n}{(\mathcal{R}_{g,2n})})$, then $\mathcal{S}(\tau_{0},\mathcal{N})=\upsilon^{-1}(\tau_{0})$ for any local irreducible component $\mathcal{N}$ of $\mathcal{N}_{n-4}$ at $\tau_{0}$.
Proof.
When $n\geq 5$, by Corollary 2.8, $\upsilon^{-1}(\tau_{0})$ is irreducible of dimension n4n-4. Since
$n-4=\dim{\upsilon^{-1}(\tau_{0})}\geq\dim_{({\boldsymbol{z}},\tau_{0})}{(\upsilon^{-1}(\tau_{0})\cap\mathcal{M})}\geq{n-4}$
for any irreducible component $\mathcal{M}$ of $\upsilon^{-1}(\mathcal{N})\cap\mathcal{S}_{m}$ and $({\boldsymbol{z}},\tau_{0})\in\mathcal{M}$, we have $\upsilon^{-1}(\tau_{0})\subset\mathcal{M}$ for any $\mathcal{M}\in\Lambda(\mathcal{N})$, hence $\upsilon^{-1}(\tau_{0})\subset\mathcal{S}(\tau_{0},\mathcal{N})$. When $n=4$, by Lemma 4.1, for $\mathcal{M}\in\Lambda(\mathcal{N})$ and ${\boldsymbol{i}},{\boldsymbol{j}}\in{I}$, there exists $\mathcal{M}_{{\boldsymbol{i}},{\boldsymbol{j}}}\in\Lambda(\mathcal{N})$ such that
$\mathcal{M}_{{\boldsymbol{i}},{\boldsymbol{j}}}=\bigl\{({\boldsymbol{z}},\tau)\in\mathbb{C}^{d}\times\mathfrak{H}_{d}\mid({\boldsymbol{z}}-\tau\begin{pmatrix}{\boldsymbol{0}}\\ {\boldsymbol{i}}\end{pmatrix}-\Delta\begin{pmatrix}{\boldsymbol{0}}\\ {\boldsymbol{j}}\end{pmatrix},\tau)\in\mathcal{M}\bigr\}.$
By Corollary 2.8, we have $\upsilon^{-1}(\tau_{0})\subset\mathcal{S}(\tau_{0},\mathcal{N})$. ∎
We define $\mathcal{R}_{1,2n}^{h}=\mathcal{R}_{1,2n}$, and for $g\geq 2$,
$\mathcal{R}_{g,2n}^{h}=\{[\tilde{C}\overset{\phi}{\rightarrow}C]\in\mathcal{R}_{g,2h}\mid\text{$C$ is a hyperelliptic curve}\}.$
Lemma 4.5.
Let $r=\frac{d(d+1)}{2}-3g-2n+3$. For $\tau_{0}\in\pi^{-1}(\operatorname{Prym}_{g,2n}{(\mathcal{R}_{g,2n}^{h})})$, there exist
$({\boldsymbol{a}}_{1},\tau_{0}),\dots,({\boldsymbol{a}}_{r},\tau_{0})\in\upsilon^{-1}(\tau_{0})=S(A_{\tau_{0}},\mathcal{L}_{\tau_{0}})$
and
$f_{1}({\boldsymbol{z}}),\dots,f_{r}({\boldsymbol{z}})\in\bigoplus_{{\boldsymbol{i}}\in{I}}\mathbb{C}\,\theta_{\boldsymbol{i}}({\boldsymbol{z}},\tau_{0})=H^{0}(A_{\tau_{0}},\mathcal{L}_{\tau_{0}})\quad$
such that quadratics
$\sum_{k=1}^{d}\sum_{l=1}^{d}\frac{\partial^{2}f_{1}}{\partial{z_{k}}\partial{z_{l}}}({\boldsymbol{a}}_{1})z_{k}z_{l},\,\dots,\,\sum_{k=1}^{d}\sum_{l=1}^{d}\frac{\partial^{2}f_{r}}{\partial{z_{k}}\partial{z_{l}}}({\boldsymbol{a}}_{r})z_{k}z_{l}\in\mathbb{C}[z_{1},\dots,z_{d}]$
are linearly independent.
Proof.
Since $r=\dim{\operatorname{Ker}{(\rho_{2})}}$, by Corollary 3.3, there exist
$({\boldsymbol{a}}_{1},\tau_{0}),\dots,({\boldsymbol{a}}_{r},\tau_{0})\in\upsilon^{-1}(\tau_{0})=S(A_{\tau_{0}},\mathcal{L}_{\tau_{0}})$
and
$\Theta_{1},\dots,\Theta_{r}\in|\mathcal{L}_{\tau_{0}}|$
such that $\operatorname{mult}_{({\boldsymbol{a}}_{j},\tau_{0})}{\Theta_{j}}=2$, and the defining quadratics of the tangent cones $\mathcal{T}_{\Theta_{j},({\boldsymbol{a}}_{j},\tau_{0})}$ for $1\leq{j}\leq{r}$ form a basis of $\operatorname{Ker}{(\rho_{2})}$. Let
$f_{j}({\boldsymbol{z}})\in\bigoplus_{{\boldsymbol{i}}\in{I}}\mathbb{C}\,\theta_{\boldsymbol{i}}({\boldsymbol{z}},\tau_{0})=H^{0}(A_{\tau_{0}},\mathcal{L}_{\tau_{0}})$
be a defining section of $\Theta_{j}\subset{A_{\tau_{0}}}$. Then the defining quadratic of the tangent cone $\mathcal{T}_{\Theta_{j},({\boldsymbol{a}}_{j},\tau_{0})}$ is given by $\sum_{k=1}^{d}\sum_{l=1}^{d}\frac{\partial^{2}f_{j}}{\partial{z_{k}}\partial{z_{l}}}({\boldsymbol{a}}_{r})z_{k}z_{l}\in\mathbb{C}[z_{1},\dots,z_{d}]$. ∎
Lemma 4.6 (Heat equation, [2] Proposition 8.5.5.).
For ${\boldsymbol{i}}\in{I}$ and $1\leq{k}\leq{l}\leq{d}$,
$\frac{\partial^{2}\theta_{\boldsymbol{i}}}{\partial{{z}_{k}}\partial{z_{l}}}({\boldsymbol{z}},\tau)=2\pi\sqrt{-1}(1+\delta_{kl})\frac{\partial\theta_{\boldsymbol{i}}}{\partial{\tau_{kl}}}({\boldsymbol{z}},\tau).$
Proof of Theorem 1.4.
Let $\mathcal{N}$ be a local irreducible component of $\mathcal{N}_{n-4}$ at $\tau_{0}\in\pi^{-1}(\operatorname{Prym}_{g,2n}{(\mathcal{R}_{g,2n}^{h})})$. By Lemma 4.5 and Lemma 4.6, there exist $({\boldsymbol{a}}_{j},\tau_{0})\in\upsilon^{-1}(\tau_{0})$ and $f_{j}({\boldsymbol{z}})\in\bigoplus_{{\boldsymbol{i}}\in{I}}\mathbb{C}\,\theta_{\boldsymbol{i}}({\boldsymbol{z}},\tau_{0})$ for $1\leq{j}\leq{r}$ such that the following quadratics are linearly independent:
$\sum_{1\leq{k}\leq{l}\leq{d}}\frac{\partial{f_{1}}}{\partial{\tau_{kl}}}({\boldsymbol{a}}_{1})z_{k}z_{l},\dots,\sum_{1\leq{k}\leq{l}\leq{d}}\frac{\partial{f_{r}}}{\partial{\tau_{kl}}}({\boldsymbol{a}}_{r})z_{k}z_{l}\in\mathbb{C}[z_{1},\dots,z_{d}].$
Hence, the cotangent vectors
$\sum_{1\leq{k}\leq{l}\leq{d}}\frac{\partial{f_{1}}}{\partial{\tau_{kl}}}({\boldsymbol{a}}_{1})d\tau_{kl},\dots,\sum_{1\leq{k}\leq{l}\leq{d}}\frac{\partial{f_{r}}}{\partial{\tau_{kl}}}({\boldsymbol{a}}_{r})d\tau_{kl}\in\bigoplus_{1\leq{k}\leq{l}\leq{d}}\mathbb{C}d\tau_{kl}$
are linearly independent. By Lemma 4.4 and Proposition 4.3, we have
$\dim{\mathcal{N}}\leq\dim{\mathfrak{H}_{d}}-r=3g+2n-3.$
Since $\mathcal{R}_{g,2n}$ is irreducible by [3, p138], there exist an irreducible component $\mathcal{P}$ of the generalized Andreotti-Mayer locus $\mathcal{N}_{d,n-4}^{\Delta}\subset\mathcal{A}_{d}^{\Delta}$ such that $\operatorname{Prym}_{g,2n}{(\mathcal{R}_{g,2n})}\subset\mathcal{P}$. Then, we have $\pi^{-1}(\mathcal{P})\subset\mathcal{N}_{n-4}$, and
$\dim{\mathcal{P}}=\dim_{\tau_{0}}{\pi^{-1}(\mathcal{P})}\leq\dim_{\tau_{0}}{\mathcal{N}}\leq 3g+2n-3,$
for some local irreducible component $\mathcal{N}$ of $\mathcal{N}_{n-4}$ at $\tau_{0}\in\pi^{-1}(\operatorname{Prym}_{g,2n}{(\mathcal{R}_{g,2n}^{h})})$. On the other hand, by Theorem 1.1, we have
$\dim{\mathcal{P}}\geq\dim{\operatorname{Prym}_{g,2n}{(\mathcal{R}_{g,2n})}}=3g+2n-3.$
Hence, we conclude that $\mathcal{P}=\overline{\operatorname{Prym}_{g,2n}{(\mathcal{R}_{g,2n})}}$. ∎